OpenStudy (anonymous):

if one root of ax^2 + bx + c =0 is doubled and another is halfved then the new quadratic equation formed has equal roots then the discriminant of ax^2 + bx + c = 0 is

OpenStudy (anonymous):

let roots = s and t 2s + t/2 = -b/a st = c/a

OpenStudy (anonymous):

is my approach correct

OpenStudy (experimentx):

seems so .. try working ahead.

OpenStudy (experimentx):

hmm .. the new equation has equal roots. don;t get confused with the original equation.

OpenStudy (anonymous):

yes!

OpenStudy (experimentx):